1) ChemCo is a price-taker able to receive $10 for each bottle of vitamins they sell....

1)

ChemCo is a price-taker able to receive $10 for each bottle of vitamins they sell. At the profit maximizing level of production their AVC = $5 and AFC = $3. Given the following points on their MC curve, what is their maximum possible profit?

Quantity 0 100 200 300 400 500 600
Marginal Cost n/a $8 $4 $6 $8 $10 $15

        

Group of answer choices

a No profit

b $4000

c $5000

d $1000

e $500

2)

Assume a price-taking firm faces the following marginal costs:

Quantity 0 1 2 3 4 5 6
Marginal Cost n/a 4 3 5 6 8 10

         

If the market price is $8 and this firm has no fixed costs, how much profit will it make?

Group of answer choices

a profit = 64

b profit = 14

c profit = 10

d profit = 40
Answer #1

1. Ans: d) $1000

Explanation:

Under perfect competition , the profit maximization condition is where price equals marginal cost ( P = MC).

So , at the profit maximizing level ;

Output ( Q ) = 500

Total Revenue = Price * Quantity = $10 * 500 = $5000

ATC = AFC + AVC = $3 + $5 = $8

Total cost = ATC * Q = $8 * 500 = $4000

Profit = Total Revenue - Total cost = $5000 - $4000 = $1000

Quantity Marginal Cost
0 n/a
100 8
200 4
300 6
400 8
500 10
600 15

2.Ans: b) profit = 14

Explanation:

Under perfect competition , the profit maximization condition is where price equals marginal cost ( P = MC).

So , at the profit maximizing level of output is 5.

Marginal cost = Change in total cost / Change in Quantity

Quantity Marginal Cost TC TR Profit / Loss
0 n/a 0 0 0
1 4 4 8 4
2 3 7 16 9
3 5 12 24 12
4 6 18 32 14
5 8 26 40 14
6 10 36 48 12